A carpenter charges a set fee for each job in addition to his hourly charge. If a carpenter bills his customer $125 for a job which takes him 3 hours and bills a second customer $200 for a 6 hour job, how much would he charge for a job that takes 8 hours?
Answer by ankor@dixie-net.com(12701)

A carpenter charges a set fee for each job in addition to his hourly charge.
If a carpenter bills his customer $125 for a job which takes him 3 hours
and bills a second customer $200 for a 6 hour job,
:
Let f = fixed fee
Let h = hourly rate
:
Two equations:
3h + f = 125; "$125 for a job which takes him 3 hours"
6h + f = 200; "$200 for a 6 hr job"
:
Subtract the 1st equation from the 2nd equation
6h + f = 200
3h + f = 125
--------------- eliminates f, find h
3h = 75
h = $25 an hr
;
Find f in the 1st equation
3(25) + f = 125
f = 125 - 75
f = $50 set fee
:
:
Check solution in the 2nd equation
6(25) + 50 = 200
150 + 50 = 200
:
how much would he charge for a job that takes 8 hours?
8(25) + 50 = $250